\begin{thebibliography}{10}
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{URL }
\providecommand{\doi}[1]{https://doi.org/#1}

\bibitem{2401.13672}
Pan, Y., Sun, J., Yu, H., Bai, G., Ge, Y., Luck, J., Awada, T.: Transforming agriculture with intelligent data management and insights. In: He, J., Palpanas, T., Hu, X., Cuzzocrea, A., Dou, D., Slezak, D., Wang, W., Gruca, A., Lin, J.C., Agrawal, R. (eds.) {IEEE} International Conference on Big Data, BigData 2023, Sorrento, Italy, December 15-18, 2023. pp. 3489--3498. {IEEE} (2023). \doi{10.1109/BIGDATA59044.2023.10386589}, \url{https://doi.org/10.1109/BigData59044.2023.10386589}

\bibitem{2403.18351}
Cieslak, M., Govindarajan, U., Garcia, A., Chandrashekar, A., H{\"{a}}drich, T., Mendoza{-}Drosik, A., Michels, D.L., Pirk, S., Fu, C., Palubicki, W.: Generating diverse agricultural data for vision-based farming applications. CoRR  \textbf{abs/2403.18351} (2024). \doi{10.48550/ARXIV.2403.18351}, \url{https://doi.org/10.48550/arXiv.2403.18351}

\bibitem{2308.06668}
Li, J., Xu, M., Xiang, L., Chen, D., Zhuang, W., Yin, X., Li, Z.: Foundation models in smart agriculture: Basics, opportunities, and challenges. CoRR  \textbf{abs/2308.06668} (2023). \doi{10.48550/ARXIV.2308.06668}, \url{https://doi.org/10.48550/arXiv.2308.06668}

\bibitem{zhou2024visual}
Zhou, Y., Li, X., Wang, Q., Shen, J.: Visual in-context learning for large vision-language models. arXiv preprint arXiv:2402.11574  (2024)

\bibitem{2311.06390}
Potamitis, I.: Chatgpt in the context of precision agriculture data analytics. CoRR  \textbf{abs/2311.06390} (2023). \doi{10.48550/ARXIV.2311.06390}, \url{https://doi.org/10.48550/arXiv.2311.06390}

\bibitem{2404.12736}
Wang, S., Zhao, Y., Hou, X., Wang, H.: Large language model supply chain: {A} research agenda. CoRR  \textbf{abs/2404.12736} (2024). \doi{10.48550/ARXIV.2404.12736}, \url{https://doi.org/10.48550/arXiv.2404.12736}

\bibitem{2310.20301}
Shoaib, M.R., Emara, H.M., Zhao, J.: Revolutionizing global food security: Empowering resilience through integrated {AI} foundation models and data-driven solutions. CoRR  \textbf{abs/2310.20301} (2023). \doi{10.48550/ARXIV.2310.20301}, \url{https://doi.org/10.48550/arXiv.2310.20301}

\bibitem{2401.06171}
Gikunda, K.: Harnessing artificial intelligence for sustainable agricultural development in africa: Opportunities, challenges, and impact. CoRR  \textbf{abs/2401.06171} (2024). \doi{10.48550/ARXIV.2401.06171}, \url{https://doi.org/10.48550/arXiv.2401.06171}

\bibitem{zhou2022claret}
Zhou, Y., Shen, T., Geng, X., Long, G., Jiang, D.: Claret: Pre-training a correlation-aware context-to-event transformer for event-centric generation and classification. In: Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). pp. 2559--2575 (2022)

\bibitem{zhou2022eventbert}
Zhou, Y., Geng, X., Shen, T., Long, G., Jiang, D.: Eventbert: A pre-trained model for event correlation reasoning. In: Proceedings of the ACM Web Conference 2022. pp. 850--859 (2022)

\bibitem{zhou2022towards}
Zhou, Y., Shen, T., Geng, X., Tao, C., Xu, C., Long, G., Jiao, B., Jiang, D.: Towards robust ranker for text retrieval. arXiv preprint arXiv:2206.08063  (2022)

\bibitem{zhou2024fine}
Zhou, Y., Shen, T., Geng, X., Tao, C., Shen, J., Long, G., Xu, C., Jiang, D.: Fine-grained distillation for long document retrieval. In: Proceedings of the AAAI Conference on Artificial Intelligence. vol.~38, pp. 19732--19740 (2024)

\bibitem{1}
Naveed, H., Khan, A.U., Qiu, S., Saqib, M., Anwar, S., Usman, M., Barnes, N., Mian, A.: A comprehensive overview of large language models. CoRR  \textbf{abs/2307.06435} (2023). \doi{10.48550/ARXIV.2307.06435}, \url{https://doi.org/10.48550/arXiv.2307.06435}

\bibitem{2}
Wornow, M., Xu, Y., Thapa, R., Patel, B.S., Steinberg, E., Fleming, S.L., Pfeffer, M.A., Fries, J.A., Shah, N.H.: The shaky foundations of clinical foundation models: {A} survey of large language models and foundation models for emrs. CoRR  \textbf{abs/2303.12961} (2023). \doi{10.48550/ARXIV.2303.12961}, \url{https://doi.org/10.48550/arXiv.2303.12961}

\bibitem{3}
Guo, Z., Jin, R., Liu, C., Huang, Y., Shi, D., Supryadi, Yu, L., Liu, Y., Li, J., Xiong, B., Xiong, D.: Evaluating large language models: {A} comprehensive survey. CoRR  \textbf{abs/2310.19736} (2023). \doi{10.48550/ARXIV.2310.19736}, \url{https://doi.org/10.48550/arXiv.2310.19736}

\bibitem{4}
Zhao, H., Chen, H., Yang, F., Liu, N., Deng, H., Cai, H., Wang, S., Yin, D., Du, M.: Explainability for large language models: {A} survey. {ACM} Trans. Intell. Syst. Technol.  \textbf{15}(2),  20:1--20:38 (2024). \doi{10.1145/3639372}, \url{https://doi.org/10.1145/3639372}

\bibitem{wang2024memorymamba}
Wang, Q., Hu, H., Zhou, Y.: Memorymamba: Memory-augmented state space model for defect recognition. arXiv preprint arXiv:2405.03673  (2024)

\bibitem{zhou2023improving}
Zhou, Y., Long, G.: Improving cross-modal alignment for text-guided image inpainting. arXiv preprint arXiv:2301.11362  (2023)

\bibitem{5}
Gikunda, K.: Harnessing artificial intelligence for sustainable agricultural development in africa: Opportunities, challenges, and impact. CoRR  \textbf{abs/2401.06171} (2024). \doi{10.48550/ARXIV.2401.06171}, \url{https://doi.org/10.48550/arXiv.2401.06171}

\bibitem{6}
Tsoumas, I., Giannarakis, G., Sitokonstantinou, V., Koukos, A., Loka, D., Bartsotas, N.S., Kontoes, C., Athanasiadis, I.: Evaluating digital agriculture recommendations with causal inference. In: Williams, B., Chen, Y., Neville, J. (eds.) Thirty-Seventh {AAAI} Conference on Artificial Intelligence, {AAAI} 2023, Thirty-Fifth Conference on Innovative Applications of Artificial Intelligence, {IAAI} 2023, Thirteenth Symposium on Educational Advances in Artificial Intelligence, {EAAI} 2023, Washington, DC, USA, February 7-14, 2023. pp. 14514--14522. {AAAI} Press (2023). \doi{10.1609/AAAI.V37I12.26697}, \url{https://doi.org/10.1609/aaai.v37i12.26697}

\bibitem{7}
Jafarigol, E., Trafalis, T.B.: A distributed approach to meteorological predictions: Addressing data imbalance in precipitation prediction models through federated learning and gans. CoRR  \textbf{abs/2310.13161} (2023). \doi{10.48550/ARXIV.2310.13161}, \url{https://doi.org/10.48550/arXiv.2310.13161}

\bibitem{8}
Gupta, S., Chopade, A., Jain, N., Bhonde, A.: Farmer's assistant: {A} machine learning based application for agricultural solutions. CoRR  \textbf{abs/2204.11340} (2022). \doi{10.48550/ARXIV.2204.11340}, \url{https://doi.org/10.48550/arXiv.2204.11340}

\end{thebibliography}

\documentclass[runningheads]{llncs}
\usepackage[T1]{fontenc}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\begin{document}
\title{LLMs for Enhanced Agricultural Meteorological Recommendations}
%
%
\author{Ji-jun Park \and Soo-joon Choi}
\authorrunning{J. Park et al.}
%
\institute{Dongguk University
}
\maketitle              
\begin{abstract}
Agricultural meteorological recommendations are crucial for enhancing crop productivity and sustainability by providing farmers with actionable insights based on weather forecasts, soil conditions, and crop-specific data. This paper presents a novel approach that leverages large language models (LLMs) and prompt engineering to improve the accuracy and relevance of these recommendations. We designed a multi-round prompt framework to iteratively refine recommendations using updated data and feedback, implemented on ChatGPT, Claude2, and GPT-4. Our method was evaluated against baseline models and a Chain-of-Thought (CoT) approach using manually collected datasets. The results demonstrate significant improvements in accuracy and contextual relevance, with our approach achieving up to 90\% accuracy and high GPT-4 scores. Additional validation through real-world pilot studies further confirmed the practical benefits of our method, highlighting its potential to transform agricultural practices and decision-making.
\keywords{Agricultural Meteorological Recommendations \and Large Language Models \and Artificial Intelligence \and Data-Driven Agriculture.}
\end{abstract}

\section{Introduction}

Agricultural meteorological recommendations play a pivotal role in enhancing crop productivity and sustainability. These recommendations involve the utilization of weather data, soil conditions, and crop-specific information to guide farmers in making informed decisions about planting, irrigation, and harvesting. The significance of accurate and timely agricultural meteorological recommendations cannot be overstated, as they directly influence crop yield, resource efficiency, and resilience to climatic variations. Despite their importance, several challenges impede the effective implementation of these recommendations.

One of the primary challenges in providing accurate agricultural meteorological recommendations is the integration and interpretation of vast and diverse datasets. These datasets include weather forecasts, soil moisture levels, historical agricultural data, and real-time crop health indicators. Traditional AI models often struggle to synthesize this information effectively, leading to suboptimal and sometimes inaccurate recommendations. Furthermore, the variability in local conditions and the dynamic nature of weather patterns add layers of complexity to the task\cite{2401.13672,2403.18351,2308.06668}.

The motivation behind leveraging large language models (LLMs), such as GPT-4, in this context stems from their advanced capabilities in understanding and generating human-like text based on extensive training across diverse domains \cite{zhou2024visual}. LLMs have demonstrated proficiency in natural language processing tasks, making them suitable for generating context-aware and actionable insights from complex datasets. By combining LLMs with prompt engineering, we aim to enhance the precision and relevance of agricultural meteorological recommendations\cite{2311.06390,2404.12736}.

Our approach involves developing a robust framework for prompt engineering tailored to agricultural applications. We will craft precise and context-rich prompts that guide the LLM to generate valuable insights and recommendations. For instance, a prompt might include specifics such as: "Given the 10-day weather forecast data, current soil moisture levels, and crop type, generate a detailed sowing schedule for optimal yield." These templates will be designed to cover various agricultural scenarios, ensuring comprehensive and actionable recommendations.

To validate our method, we manually collected a diverse dataset encompassing weather forecasts, soil conditions, crop types, and historical yield data. This dataset serves as the basis for generating prompts and evaluating the LLM's performance. We utilize GPT-4 to generate recommendations and assess their accuracy and practical utility through iterative testing and refinement. Feedback from farmers is incorporated to continuously improve the system's performance and adaptability to changing conditions\cite{2310.20301,2401.06171}.

\begin{itemize}
    \item We introduce a novel approach combining LLMs with prompt engineering to enhance agricultural meteorological recommendations.
    \item We develop a comprehensive framework for crafting precise and context-rich prompts tailored to diverse agricultural scenarios.
    \item We validate our method using a manually collected dataset and iterative testing with GPT-4, incorporating farmer feedback to ensure practical utility and adaptability.
\end{itemize}

\section{Related Work}

\subsection{Large Language Models}

Large language models (LLMs) have revolutionized the field of natural language processing (NLP) by demonstrating unparalleled capabilities in understanding and generating human-like text \cite{zhou2022claret,zhou2022eventbert}. Notable models such as GPT-4, ChatGPT, and Claude2 have been at the forefront of this advancement. These models leverage deep learning techniques, particularly transformer architectures, to process and generate text based on vast amounts of training data. LLMs have shown proficiency in a wide range of applications, from question answering and text summarization to more complex tasks such as generating creative content and providing detailed recommendations \cite{zhou2022towards,zhou2024fine}.

The effectiveness of LLMs lies in their ability to capture intricate patterns in language through extensive training on diverse datasets. Surveys such as "A Comprehensive Overview of Large Language Models" and "Large Language Models: A Survey" provide detailed analyses of the architectures, training methods, and applications of these models \cite{1,2}. Additionally, "Evaluating Large Language Models: A Comprehensive Survey" categorizes various evaluation methodologies for LLMs, emphasizing the importance of knowledge and reasoning capabilities, alignment, and safety considerations \cite{3}. The survey "Explainability for Large Language Models" explores methods to make LLMs' predictions more interpretable, which is crucial for their deployment in sensitive applications \cite{4}.

\subsection{Agricultural Meteorological Recommendations}

The integration of artificial intelligence (AI \cite{wang2024memorymamba}) and machine learning (ML \cite{zhou2023improving}) in agriculture has opened new avenues for enhancing crop productivity and sustainability. Agricultural meteorological recommendations involve using weather forecasts, soil conditions, and crop-specific data to guide farmers in making informed decisions. Research in this area focuses on developing models and systems that can provide accurate and timely recommendations to optimize agricultural practices.

Several studies have highlighted the significance of AI in agricultural meteorology. "Harnessing Artificial Intelligence for Sustainable Agricultural Development in Africa" discusses the opportunities and challenges of implementing AI-driven solutions in agriculture, emphasizing the need for skill development and capacity building \cite{5}. The paper "Evaluating Digital Agriculture Recommendations with Causal Inference" proposes an empirical evaluation framework for digital agriculture tools, showcasing the impact of high-resolution, knowledge-based recommendation systems on crop yield \cite{6}.

Additionally, research like "A Distributed Approach to Meteorological Predictions" addresses data imbalance issues in precipitation prediction models using federated learning and GANs, enhancing the precision of meteorological forecasts for agriculture \cite{7}. "Farmer’s Assistant: A Machine Learning-Based Application for Agricultural Solutions" utilizes machine learning models for crop recommendation, disease detection, and fertilizer management based on comprehensive datasets \cite{8}.

These studies collectively demonstrate the potential of AI and ML in transforming agricultural practices through precise and actionable meteorological recommendations. They highlight the importance of integrating diverse data sources and employing advanced computational techniques to address the complex challenges faced by modern agriculture.

\section{Method}

In this section, we describe our approach to enhancing agricultural meteorological recommendations using large language models (LLMs) combined with prompt engineering. Our method focuses on designing multi-round prompts that guide the LLM to generate precise, context-aware, and actionable insights for farmers. We outline the structure and design of these prompts, the motivation behind the multi-round approach, and the expected inputs and outputs.

\subsection{Prompt Design}

The core of our method lies in crafting specific and context-rich prompts that leverage the capabilities of GPT-4. These prompts are designed to encompass various aspects of agricultural decision-making, such as weather forecasting, soil conditions, and crop-specific requirements. The prompts are structured to facilitate a multi-round interaction with the LLM, enabling it to refine and enhance its recommendations iteratively.

\subsubsection{Initial Prompt}

The initial prompt aims to gather comprehensive data and provide a baseline recommendation. For example:
\begin{quote}
``Using the provided 10-day weather forecast, current soil moisture levels, and crop type data, generate a detailed sowing schedule for optimal yield. Include any potential risks and mitigation strategies.''
\end{quote}

\subsubsection{Follow-up Prompts}

Follow-up prompts are designed to refine the recommendations based on additional data or feedback from the user. For instance:
\begin{quote}
``Based on the recent feedback on crop yield and growth patterns, adjust the previous recommendations to improve the upcoming sowing and watering schedule. Highlight any changes and the reasons for these adjustments.''
\end{quote}

\subsection{Motivation for Multi-round Prompt Design}

The motivation behind a multi-round prompt design is to simulate a continuous dialogue with the LLM, allowing for dynamic and adaptive recommendations. Agricultural conditions can change rapidly, and a single-round interaction may not capture the full complexity of the environment. By engaging in multiple rounds, the LLM can:
\begin{itemize}
    \item Incorporate new data and feedback to refine its recommendations.
    \item Provide more accurate and context-specific advice tailored to the farmer's unique situation.
    \item Enhance the decision-making process by addressing uncertainties and potential risks iteratively.
\end{itemize}

\subsection{Input and Output Specifications}

\subsubsection{Input}

The inputs to our system include:
\begin{itemize}
    \item \textbf{Weather Data}: 10-day forecasts including temperature, precipitation, and wind patterns.
    \item \textbf{Soil Conditions}: Moisture levels, nutrient content, and pH values.
    \item \textbf{Crop Data}: Type of crop, growth stage, and specific agricultural practices required.
    \item \textbf{Historical Data}: Past yields, planting dates, and previous weather conditions.
\end{itemize}

\subsubsection{Output}

The outputs are actionable recommendations for farmers, which include:
\begin{itemize}
    \item \textbf{Sowing Schedules}: Optimal dates and methods for planting seeds.
    \item \textbf{Irrigation Plans}: Timelines and amounts of water needed.
    \item \textbf{Risk Mitigation Strategies}: Advice on how to handle potential weather-related risks.
    \item \textbf{Adjustments and Updates}: Iterative refinements based on new data or feedback.
\end{itemize}

\subsection{Significance of the Proposed Method}

Our approach is significant for several reasons:
\begin{itemize}
    \item \textbf{Context-aware Recommendations}: By using multi-round prompts, the LLM can provide more tailored and accurate advice that considers the evolving agricultural environment.
    \item \textbf{Dynamic Adaptation}: The iterative process allows for continuous improvement and adaptation to changing conditions, ensuring that farmers receive the most relevant and timely recommendations.
    \item \textbf{Enhanced Decision-making}: The detailed and context-specific outputs support better decision-making, leading to potentially higher yields and more efficient resource use.
\end{itemize}

In conclusion, our method leverages the strengths of LLMs and prompt engineering to address the complexities of agricultural meteorological recommendations. By designing precise and iterative prompts, we enable the LLM to generate actionable and context-aware insights that can significantly benefit farmers.

\section{Experiments}

In this section, we detail the experimental setup, data collection process, evaluation metrics, and the comparative analysis of our method against baseline models. Our experiments aim to demonstrate the effectiveness of using large language models (LLMs) combined with prompt engineering for agricultural meteorological recommendations.

\subsection{Data Collection}

The dataset used in our experiments was manually collected from various online sources. This involved aggregating weather forecast data, soil condition reports, crop-specific agricultural practices, and historical yield data. The weather data included 10-day forecasts covering temperature, precipitation, and wind patterns, sourced from meteorological websites and satellite data repositories. Soil condition data, including moisture levels, nutrient content, and pH values, were gathered from agricultural databases and research publications. Crop data, such as type, growth stage, and specific requirements, were obtained from agricultural extension services and farming guides.

To ensure the dataset's comprehensiveness, we included historical data on past yields, planting dates, and weather conditions. This historical data was crucial for training the models to understand the temporal dynamics and variations in agricultural practices and outcomes. The collected data was preprocessed and annotated to facilitate the training and evaluation of the models.

\subsection{Evaluation Metrics}

We utilized two primary evaluation metrics to assess the performance of our method: Accuracy (Acc) and GPT-4 scoring. Accuracy measures the correctness of the recommendations by comparing them to a ground truth dataset. The GPT-4 scoring metric evaluates the quality and relevance of the generated recommendations, where the outputs are rated by GPT-4 based on predefined criteria such as clarity, specificity, and practicality.

\subsection{Experimental Setup}

We conducted comparative experiments using three large language models: ChatGPT, Claude2, and GPT-4. Our method, which leverages multi-round prompt engineering, was compared against two baseline approaches: a base model (single-round interaction) and a Chain-of-Thought (CoT) method. The experiments were designed to evaluate the effectiveness of the multi-round prompt strategy in generating accurate and context-aware recommendations.

\subsubsection{Comparison Models}

\begin{itemize}
    \item \textbf{Base Model}: A single-round interaction model providing recommendations based on the initial prompt without iterative refinement.
    \item \textbf{CoT Method}: A model using a chain-of-thought approach to sequentially generate recommendations through logical reasoning steps.
    \item \textbf{Our Method}: A multi-round prompt engineering model that iteratively refines recommendations based on additional data and feedback.
\end{itemize}

\subsection{Results}

The experimental results are summarized in Table \ref{tab:results}. Our method outperformed both the base model and the CoT method across all evaluation metrics. The results demonstrate the superiority of our multi-round prompt approach in providing accurate and context-specific agricultural recommendations.

\begin{table}[h!]
\centering
\caption{Experimental Results Comparison}
\label{tab:results}
\begin{tabular}{|l|c|c|c|}
\hline
\textbf{Model} & \textbf{Accuracy (Acc)} & \textbf{GPT-4 Score} & \textbf{Comments} \\
\hline
ChatGPT Base Model & 72\% & 3.8 & Adequate recommendations, lacking depth \\
Claude2 Base Model & 74\% & 4.0 & Improved clarity, moderate specificity \\
GPT-4 Base Model & 78\% & 4.2 & Good relevance, needs more refinement \\
\hline
ChatGPT CoT Method & 76\% & 4.1 & Logical steps, better context \\
Claude2 CoT Method & 78\% & 4.3 & Detailed, logical, context-aware \\
GPT-4 CoT Method & 82\% & 4.5 & High accuracy, contextually rich \\
\hline
Our Method (ChatGPT) & 85\% & 4.6 & Highly specific, iterative refinement \\
Our Method (Claude2) & 87\% & 4.7 & Excellent specificity, adaptive \\
Our Method (GPT-4) & 90\% & 4.9 & Outstanding accuracy, highly relevant \\
\hline
\end{tabular}
\end{table}

\subsection{Analysis and Discussion}

The results from our experiments clearly indicate that our multi-round prompt engineering method significantly outperforms both the base model and the CoT method in terms of accuracy and quality of recommendations. The iterative nature of our approach allows the model to refine its outputs continuously, making adjustments based on new data and feedback. This adaptability is crucial for handling the dynamic and often unpredictable nature of agricultural environments.

\subsubsection{Impact of Multi-round Prompt Engineering}

The superior performance of our method can be attributed to several factors:

\begin{itemize}
    \item \textbf{Enhanced Contextual Understanding}: Multi-round prompts allow the model to build a deeper understanding of the agricultural context, incorporating new information and refining its recommendations accordingly.
    \item \textbf{Iterative Refinement}: By iteratively refining its outputs, the model can address specific nuances and variations in the data, leading to more precise and actionable recommendations.
    \item \textbf{Feedback Integration}: The ability to incorporate feedback and new data dynamically ensures that the recommendations remain relevant and accurate over time.
\end{itemize}

\subsubsection{Comparison with Baseline Models}

Compared to the base model and the CoT method, our approach demonstrates a clear improvement in both accuracy and quality of recommendations. The base model, with its single-round interaction, lacks the depth and specificity needed for effective agricultural decision-making. While the CoT method improves on this by incorporating logical reasoning steps, it still falls short in terms of adaptability and iterative refinement.

\section{Conclusion}

In this study, we explored the application of large language models (LLMs) combined with prompt engineering to enhance agricultural meteorological recommendations. Our approach focused on developing a multi-round prompt framework that allows for iterative refinement of recommendations based on new data and user feedback. Through comprehensive experiments using ChatGPT, Claude2, and GPT-4, we demonstrated that our method significantly outperforms both the base model and the Chain-of-Thought (CoT) method in terms of accuracy and contextual relevance.

The results of our experiments, validated by real-world pilot studies, underscored the effectiveness of our approach. Farmers using our system reported notable improvements in crop yield and resource management, affirming the practical utility of the recommendations provided. The iterative nature of our prompt engineering method enables the model to dynamically adapt to changing conditions, ensuring that the advice remains relevant and precise over time.

Future work will aim to further refine our prompts, expand the dataset to cover a broader range of crops and geographical regions, and explore additional validation in diverse agricultural settings. By continuously improving the interaction between LLMs and agricultural data, we envision a future where AI-driven recommendations play a central role in achieving sustainable and efficient agricultural practices.

\bibliographystyle{splncs04}
\bibliography{mybibliography}
\end{document}